# Surfaces and other Peano Continua with no Generic Chains

### v1

@@ 7. Classifiability and Dynamics
Recall that a Polish group has the generic point property if every minimal flow has a comeager orbit. Given a Peano continuum $X$ and a subgroup $G\leq\operatorname{Homeo}(X)$, we can obtain the failure of the generic point property for $G$ directly from Theorem 1.2, if we can prove that the action $G\curvearrowright\Phi(X)$ is minimal. The first thing to notice is that a necessary prerequisite for this is the minimality the action $G\curvearrowright X$, since we recall that the map from $\Phi(X)$ onto $X$ associating to each chain $\mathcal{C}$ its root $\bigcap\mathcal{C}$ is a continuous $\operatorname{Homeo}(X)$-equivariant map. Therefore, for any compact manifold with boundary, the action on the space of chains is not minimal. Minimality of $G\curvearrowright X$ is not sufficient, though: while $\operatorname{Homeo}(S^{1})$ acts minimally, even transitively on the circle $S^{1}$, $\operatorname{Homeo}(S^{1})\curvearrowright\Phi(X)$ has a proper closed invariant subset: take the set of those chains $\mathcal{C}$ such that for each $C\in\mathcal{C}\setminus\{S^{1}\}$ the root of $\mathcal{C}$ belongs to the boundary of $C$.
@@
Let $X$ be a Peano continuum with no locally separating points. Let $G\leq\operatorname{Homeo}(X)$ be such that $G\curvearrowright X$ is minimal. Suppose that for each $\mathcal{C}\in\Phi(X)$ with root $x_{0}$, and any connected open sets $U_{0},U_{1},V$ with $x_{0}\in U_{0}$ and $U_{0},U_{1}\subseteq V$, there are $C\in\mathcal{C}$ and $g\in G$ such that:
@@
The proof relies on the following lemma, whose straightforward proof is left to the reader.
@@
Let $\mathcal{C}\in\Phi(X)$ with root $x_{0}$ and a nonempty open $\mathbf{U}\subseteq\Phi(X)$ be given. By Proposition 5.2 and (Upgrading Walks to Paths). (Theorem 6.2), there are a brick partition $\mathcal{U}$ of $X$ and a spaced path $u$ on $\mathcal{U}$ such that $\mathbf{O}_{\mathcal{U},u}\subseteq\mathbf{U}$.
@@
We briefly recall what it means for an equivalence relation on a Polish space to be classifiable by countable structures and refer the reader to [hjorth2000classification]*Section 2.3 for more details. Given a countable relational language =$\mathcal{L}=\{R_{i}\}_{i\in I}$, where each $R_{i}$ has arity $n_{i}$, we topologize the collection $\mathrm{Mod}(\mathcal{L})$ of all $\mathcal{L}$-structures with domain $\mathbb{N}$, by identifying it with the product space
@@
where every $R_{i}\subseteq\mathbb{N}^{n(i)}$ is identified with its indicator function. Consider now another Polish space $Y$ and an equivalence relation $\equiv$ on $Y$. We say that $(Y,\equiv)$ is classifiable by countable structures if and only if there exists a Borel function $f\colon Y\to\mathrm{Mod}(\mathcal{L})$, for some countable relational language $\mathcal{L}$, such that for all $y,y^{\prime}\in Y$,
(7.1) $y\equiv y^{\prime}\iff f(y)\approx f(y^{\prime}).$
In order to prove Corollary 1.8, we rely on the notion of turbulence and on a classic theorem due to Hjorth.
Definition 7.3 ([hjorth2000classification]).
Let $G\curvearrowright Y$ be a continuous action on a topological space. Given a point $y\in Y$, an open neighborhood $\mathbf{U}$ of $y$, and an open neighborhood $G_{0}$ of the identity in $G$, the $(\mathbf{U},G_{0})$-local orbit of $y$ is the set of all $y^{\prime}\in\mathbf{U}$ such that there is a finite sequence $y=y_{0},\dots,y_{n}=y^{\prime}$ of elements of $\mathbf{U}$ verifying $y_{i+1}\in G_{0}\cdot y_{i}$ for all $i<n$. We shall say that all local orbits are somewhere dense if for any $y,\mathbf{U},G_{0}$ as above, the closure of the $(\mathbf{U},G_{0})$-local orbit of $y$ has nonempty interior.
An action is turbulent if it is minimal, it has no comeager orbit and all local orbits are somewhere dense.
Theorem 7.4 ([hjorth2000classification]*Corollary 3.19).
Let $G$ be a Polish group and $G\curvearrowright Y$ be a continuous action on a Polish space. If the action is turbulent then the equivalence relation $y\equiv_{G}y^{\prime}$ iff $y^{\prime}\in G\cdot y$ is not classifiable by countable structures.
@@ 7.1. Strongly Locally Homogeneous Spaces
In Theorem 7.7 below we show that the action $G\curvearrowright\Phi(X)$ is minimal and all local orbits are somewhere dense, as long as $X$ satisfies a strong form of homogeneity and $G\leq\operatorname{Homeo}(X)$ is large enough. We begin by making these last two conditions precise.
@@
Spaces $X$ such that $\operatorname{Homeo}(X)\curvearrowright X$ is locally transitive are called strongly locally homogeneous (see for instance [MR2674033]). All closed manifolds, as well as the Hilbert cube, are strongly locally homogeneous. Furthermore, $\operatorname{Homeo}_{0}(X)$ acts locally transitively on $X$ whenever $X$ is a closed manifold, as do $\operatorname{Homeo}_{+}(X)$ and $\mathrm{Diffeo}(X)$ when $X$ is moreover orientable or smooth, respectively [gutman_minimal]*Examples 3.1. Other examples of strongly locally homogeneous Peano continua are the Menger curve $\mu^{1}$ and the universal $k$-dimensional compactum $\mu^{k}$, for $k>1$ [MR0920964]*Theorem 3.2.2.
@@
As a consequence, strongly locally homogeneous Peano continuum with no locally separating points are homogeneous. Any Peano continuum which is homogeneous and is not the circle or a closed surface, is locally non-planar [MR2674033]*Remark 8. It then follows from Theorem 6.4 and Theorem 6.6 that:
@@
We now turn to the main theorem of this section. We remark that minimality of $G\curvearrowright\Phi(X)$, for $X$ a closed manifold of dimension at least 2 and $G\leq\operatorname{Homeo}(X)$ a subgroup which acts locally transitively on $X$, was already proved in [gutman_minimal]*Theorem 6.5.
@@
Let $X$ be a Peano continuum with no locally separating points. Let $G\leq\operatorname{Homeo}(X)$ be such that $G\curvearrowright X$ is locally transitive. Then the action $G\curvearrowright\Phi(X)$ is minimal and all local orbits are somewhere dense.
@@
By Proposition 7.5, $G\curvearrowright X$ is minimal, so to prove minimality of $G\curvearrowright\Phi(X)$ it is enough to check that the conditions of Proposition 7.1 are satisfied. To this end, let $\mathcal{C}\in\Phi(X)$ with root $x_{0}$ be given, together with open connected sets $U_{0},U_{1},V$ such that $x_{0}\in U_{0}$ and $U_{0},U_{1}\subseteq V$.
Let $C\in\mathcal{C}$ be non-trivial and such that $C\subseteq V$, and pick $r_{0}\in C\setminus\{x_{0}\}$. Fix moreover $r_{1}\in U_{1}$. By Proposition 7.5, there is $g\in G_{V}$ such that $g(x_{0})=x_{0}\in U_{0}$ and $g(r_{0})=r_{1}\in U_{1}\cap g[C]$. As $g$ is the identity outside $V$, we get $g[C]\subseteq V$, hence the action $G\curvearrowright\Phi(X)$ is minimal by Proposition 7.1.
Towards proving that all local orbits are somewhere dense, fix $\mathcal{C}\in\Phi(X)$ with root $x_{0}$, an open neighborhood $\mathbf{U}$ of $\mathcal{C}$ and an open neighborhood $G_{0}$ of the identity in $G$. By Proposition 5.2 and (Upgrading Walks to Paths). (Theorem 6.2), there are a brick partition $\mathcal{U}$ and a spaced path $u$ on $\mathcal{U}$ such that $\mathbf{O}_{\mathcal{U},u}\subseteq\mathbf{U}$.
We prove that the closure of the $(\mathbf{U},G_{0})$-local orbit of $\mathcal{C}$ contains $\mathbf{O}_{\mathcal{U},u}$. To this end, let $\mathbf{V}\subseteq\mathbf{O}_{\mathcal{U},u}$ be open and nonempty. By Proposition 5.2, there are a brick partition and a walk $(\mathcal{V},v)\preceq(\mathcal{U},u)$ such that $\mathbf{O}_{\mathcal{V},v}\subseteq\mathbf{V}$. Up to a refinement, by (Upgrading Walks to Paths)., we can suppose that $v$ is a spaced path, and that $\mathcal{V}$ is fine enough that
@@
By picking $\mathcal{V}$ fine enough we can moreover assume there is $V_{0}\in\mathcal{V}(u(0))$ such that $x_{0}\in\operatorname{Cl}(V_{0})$ and such that $[V_{0}]_{1}\subseteq u(0)$. Fix a path $P$ from $V_{0}$ to $v(0)$ in $\mathcal{V}(u(0))$. We will move $x_{0}$ along $P$, step by step. As a preliminary move, let $\mathcal{C}_{-1}\coloneqq\mathcal{C}$ and, using Proposition 7.5, find $g_{-1}\in G_{[V_{0}]_{1}}\subseteq G_{\mathcal{V},1}$ mapping $x_{0}$ to some $y_{0}\in V_{0}=P(0)$ and set $\mathcal{C}_{0}\coloneqq g_{-1}\cdot\mathcal{C}$. Note that $\mathcal{C}_{0}\in\mathbf{O}_{\mathcal{U},u}$ since $g_{-1}$ fixes every element of $u$ setwise. Suppose next we are given $\mathcal{C}_{i}\in\mathbf{O}_{\mathcal{U},u}$ with root $y_{i}\in P(i)$, for some $i<\operatorname{lh}(P)-1$. By Proposition 7.5, there is $g_{i}\in G_{P(i)\vee P(i+1)}$ which maps $y_{i}$ to some point $y_{i+1}\in P(i+1)$. Again, since $g_{i}$ fixes every element of $u$ setwise, $\mathcal{C}_{i+1}\coloneqq g_{i}\cdot\mathcal{C}_{i}$ belongs to $\mathbf{O}_{\mathcal{U},u}$.
@@
Suppose that $\mathcal{D}_{j}\in\mathbf{O}_{\mathcal{U},u}$ with root $z_{0}$ is given, and $\mathcal{D}_{j}\in\mathbf{O}_{\mathcal{V},\langle v(0),\dots,v(j)\rangle}$ for some $11\leq j<\operatorname{lh}(v)-$ 1. Let $D_{j}\in\mathcal{D}_{j}$ be such that $D_{j}\in O(v(0),v(1),\dots,v(j))$, and let $t_{j}\in v(j)\cap D_{j}$. In case $j=0$, make sure that $t_{0}\not=z_{0}$. By Proposition 7.5, there is $h_{j}\in G_{v(j)\vee v(j+1)}$ which maps $t_{j}$ to some point $t_{j+1}\in v(j+1)$. Again, if $j=0$, pick $h_{j}$ so that it also satisfies $h_{0}(z_{0})=z_{0}$. Call $\mathcal{D}_{j+1}\coloneqq h_{j}\cdot\mathcal{D}_{j}$. Notice that the root of $\mathcal{D}_{j+1}$ is $h_{j}(z_{0})=z_{0}$, since for $j>0$ the set $v(0)$ is pointwise fixed by $h_{j}$. Moreover, $h_{j}[D_{j}]\subseteq\bigvee_{k\leq j+1}v(k)$, as the set on the right is setwise fixed by $h_{j}$. Lemma 7.2 then implies
@@
It is left to prove that $\mathcal{D}_{j+1}\in\mathbf{O}_{\mathcal{U},u}$. If $\rho^{\mathcal{V}}_{\mathcal{U}}\cdot\langle v(0),\dots,v(j+1)\rangle\sqsubseteq u$, then $\mathbf{O}_{\mathcal{V},\langle v(0),\dots,v(j+1)\rangle}\subseteq\mathbf{O}_{\mathcal{U},u}$, in which case we conclude by (7.3). Otherwise, it holds that $v(j)\vee v(j+1)\subseteq\bigvee_{k<\operatorname{lh}(u)}u(k)$. By inductive assumption $\mathcal{D}_{j}\in\mathbf{O}_{\mathcal{U},u}$, so there is $A\in\mathcal{D}_{j}$ with $A\in O(u(0),\dots,u(-1))$. Therefore $A\cap u(-1)\neq\emptyset$, and since $h_{j}$ fixes $u(-1)$, we can infer $h_{j}[A]\cap u(-1)\neq\emptyset$. Moreover, $h_{j}[A]\subseteq\bigvee_{k<\operatorname{lh}(u)}u(k)$ since the set on the right is setwise fixed by $h_{j}$. We conclude by Lemma 7.2.
@@
Let $X$ be a strongly locally homogeneous Peano continuum which is not the circle, the sphere, or the real projective plane. Then $\operatorname{Homeo}(X)$ does not have the generic point property. In particular, its universal minimal flow is not metrizable.

### v3

@@ 7. Classifiability and Dynamics
Recall that a Polish group has the generic point property if every minimal flow has a comeager orbit. Given a Peano continuum $X$ and a subgroup $G\leq\operatorname{Homeo}(X)$, we can obtain the failure of the generic point property for $G$ directly from Theorem 1.2, if we can prove that the action $G\curvearrowright\Phi(X)$ is minimal. The first thing to notice is that a necessary prerequisite for this is the minimality the action $G\curvearrowright X$, since we recall that the map from $\Phi(X)$ onto $X$ associating to each chain $\mathcal{C}$ its root $\bigcap\mathcal{C}$ is a continuous $\operatorname{Homeo}(X)$-equivariant map. Therefore, for any compact manifold with boundary, the action of the homeomorphism group on the space of chains is not minimal. Minimality of $G\curvearrowright X$ is not sufficient, though: while $\operatorname{Homeo}(S^{1})$ acts minimally, even transitively on the circle $S^{1}$, $\operatorname{Homeo}(S^{1})\curvearrowright\Phi(X)$ has a proper closed invariant subset: take the set of those chains $\mathcal{C}$ such that for each $C\in\mathcal{C}\setminus\{S^{1}\}$ the root of $\mathcal{C}$ belongs to the boundary of $C$.
@@
Let $X$ be a Peano continuum with no locally separating points. Let $G\leq\operatorname{Homeo}(X)$ be such that $G\curvearrowright X$ is minimal. Suppose that for each $\mathcal{C}\in\Phi(X)$ with root $x_{0}$, and any connected nonempty open sets $U_{0},U_{1},V$ with $U_{0},U_{1}\subseteq V$, there are $C\in\mathcal{C}$ and $g\in G$ such that:
@@
The proof relies on the following straightforward lemma.
@@
Proof.
Since $u$ is spaced, $x_{0}\in C$, and $C\in O(u(0),\dots,u(-1))$, it holds that any subcontinuum $C^{\prime}$ of $C$ which contains $x_{0}$ is such that
[…]
By maximality and connectedness of $\mathcal{C}$, for each $i<\operatorname{lh}(u)-1$, there is $x_{0}\in C_{i}\subseteq C$ in $\mathcal{C}$ such that $C_{i}\cap u(i)\neq\emptyset$ but $C_{i}\cap u(i+1)=\emptyset$. By (7.1), $C_{i}\in O(u(0),\dots,u(i))$, and we are done. ∎
@@
Let $\mathcal{C}\in\Phi(X)$ with root $x_{0}$ and a nonempty open $\mathbf{U}\subseteq\Phi(X)$ be given. By Corollary 5.4 and (Upgrading Walks to Paths). (Theorem 6.2), there are a brick partition $\mathcal{U}$ of $X$ and a spaced path $u$ on $\mathcal{U}$ such that $\mathbf{O}_{\mathcal{U},u}\subseteq\mathbf{U}$.
@@
We briefly recall what it means for an equivalence relation on a Polish space to be classifiable by countable structures and refer the reader to [MR1967835] or [hjorth2000classification] for more details. Given a countable relational language =$\mathcal{L}=\{R_{i}\}_{i\in I}$, where each $R_{i}$ has arity $n_{i}$, we topologize the collection $\mathrm{Mod}(\mathcal{L})$ of all $\mathcal{L}$-structures with domain $\mathbb{N}$, by identifying it with the product space
@@
where every $R_{i}\subseteq\mathbb{N}^{n(i)}$ is identified with its indicator function. Consider now another Polish space $Y$ and an equivalence relation $\equiv$ on $Y$. We say that $(Y,\equiv)$ is classifiable by countable structures if there exists a Borel function $f\colon Y\to\mathrm{Mod}(\mathcal{L})$, for some countable relational language $\mathcal{L}$, such that for all $y,y^{\prime}\in Y$,
(7.2) $y\equiv y^{\prime}\iff f(y)\cong f(y^{\prime}).$
In order to prove Corollary 1.8, we rely on the notion of turbulence and on a classic theorem due to Hjorth [hjorth2000classification]. The following definitions and statements can be found in [MR1967835]*Sections 7 and 8.
Definition 7.3.
Let $G\curvearrowright Y$ be a continuous action on a topological space. Given a point $y\in Y$, an open neighborhood $\mathbf{U}$ of $y$, and an open neighborhood $G_{0}$ of the identity in $G$, the $(\mathbf{U},G_{0})$-local orbit of $y$ is the set of all $y^{\prime}\in\mathbf{U}$ such that there is a finite sequence $y=y_{0},\dots,y_{n}=y^{\prime}$ of elements of $\mathbf{U}$ verifying $y_{i+1}\in G_{0}\cdot y_{i}$ for all $i<n$. We shall say that a point $y\in Y$ is turbulent if for any $\mathbf{U},G_{0}$ as above, the closure of the $(\mathbf{U},G_{0})$-local orbit of $y$ has nonempty interior.
An action is generically turbulent if it is minimal, has no comeager orbit and comeagerly many points are turbulent, see [MR1967835]*Proposition 8.7.
Recall that the orbit equivalence relation $\equiv_{G}$ on $Y$ associated to $G\curvearrowright Y$ is defined as $y\equiv_{G}y^{\prime}$ if and only if $y^{\prime}\in G\cdot y$.
Theorem 7.4 ([hjorth2000classification]).
Let $G$ be a Polish group and $G\curvearrowright Y$ be a continuous action on a Polish space. The action is generically turbulent if and only if the orbit equivalence relation $\equiv_{G}$ is not classifiable by countable structures.
@@ 7.1. Strongly Locally Homogeneous Spaces
In Theorem 7.7 below we show that the action $G\curvearrowright\Phi(X)$ is minimal and comeagerly many point are turbulent, as long as $X$ satisfies a strong form of homogeneity and $G\leq\operatorname{Homeo}(X)$ is large enough. We begin by making these last two conditions precise.
@@
Spaces $X$ such that $\operatorname{Homeo}(X)\curvearrowright X$ is locally transitive are called strongly locally homogeneous (see for instance [MR2674033]). All closed manifolds, as well as the Hilbert cube, are strongly locally homogeneous. Furthermore, $\operatorname{Homeo}_{0}(X)$ acts locally transitively on $X$ whenever $X$ is a closed manifold, as do $\operatorname{Homeo}_{+}(X)$ and $\mathrm{Diffeo}(X)$ when $X$ is moreover orientable or smooth, respectively [gutman_minimal]*Examples 3.1. Other examples of strongly locally homogeneous Peano continua with no locally separating points are the Menger curve $\mu^{1}$ and the universal $k$-dimensional compactum $\mu^{k}$, for $k>1$ [MR0920964]*Theorem 3.2.2.
@@
As a consequence, strongly locally homogeneous Peano continuum with no locally separating points are homogeneous. Any Peano continuum which is homogeneous and is not the circle or a closed surface, is locally non-planar [MR2674033]*Remark 8. It then follows from Theorem 1.2 and Proposition 6.13 that:
@@
We now turn to the main theorem of this section. We remark that minimality of $G\curvearrowright\Phi(X)$, for $X$ a closed manifold of dimension at least 2 and $G\leq\operatorname{Homeo}(X)$ a subgroup which acts locally transitively on $X$, was already proved in [gutman_minimal]*Theorem 6.5222Gutman’s theorem applies to Peano continua which are strongly arcwise inseparable, a property that implies that points are not locally separating, but that fails for any space containing a separating arc, as is the case for the Menger curve [MR0096180]. . The following result applies to all the examples of strongly locally homogeneous Peano continua which were described at the beginning of the subsection. In particular, to the Menger curve and its group of homeomorphisms.
@@
Let $X$ be a Peano continuum with no locally separating points. Let $G\leq\operatorname{Homeo}(X)$ be such that $G\curvearrowright X$ is locally transitive. Then the action $G\curvearrowright\Phi(X)$ is minimal and comeagerly many points are turbulent.
@@
By Proposition 7.5, $G\curvearrowright X$ is minimal, so to prove minimality of $G\curvearrowright\Phi(X)$ it is enough to check that the conditions of Proposition 7.1 are satisfied. To this end, let $\mathcal{C}\in\Phi(X)$ with root $x_{0}$ be given, together with nonempty open connected sets $U_{0},U_{1},V$ such that $U_{0},U_{1}\subseteq V$.
By minimality of $G\curvearrowright X$, up to a translation, we can assume that $x_{0}\in U_{0}$. Let $C\in\mathcal{C}$ be non-trivial and such that $C\subseteq V$, and pick $x_{1}\in C\setminus\{x_{0}\}$. Fix moreover $y_{1}\in U_{1}$ distinct from $x_{0}$. By Proposition 7.5, there is $g\in G_{V}$ such that $g(x_{0})=x_{0}\in U_{0}$ and $g(x_{1})=y_{1}\in U_{1}\cap g[C]$. As $g$ is the identity outside $V$, we get $g[C]\subseteq V$, hence the action $G\curvearrowright\Phi(X)$ is minimal by Proposition 7.1.
Towards proving that comeagerly many points of $\Phi(X)$ are turbulent, fix a sequence $\mathcal{U}_{n}$ of brick partitions of $X$ such that $\mathcal{U}_{n+1}\preceq\mathcal{U}_{n}$ and the mesh of $\mathcal{U}_{n}$ tends to zero. By Proposition 5.2, the set $Z\coloneqq\bigcap_{n\in\mathbb{N}}\bigcup\Phi(\mathcal{U}_{n})$ is comeager in $\Phi(X)$. Let $\mathcal{C}\in Z$ with root $x_{0}$, an open neighborhood $\mathbf{U}$ of $\mathcal{C}$, and an open neighborhood $G_{0}$ of the identity in $G$ be given. Since $\mathcal{C}\in Z$, there are $n\in\mathbb{N}$ and a walk $u$ on $\mathcal{U}_{n}$ such that $\operatorname{Im}(u)=\mathcal{U}_{n}$ and $\mathcal{C}\in\mathbf{O}_{\mathcal{U}_{n},u}\subseteq\mathbf{U}$. To lighten notation, let us rename $\mathcal{U}_{n}$ to $\mathcal{U}$.
We prove that the closure of the $(\mathbf{U},G_{0})$-local orbit of $\mathcal{C}$ contains $\mathbf{O}_{\mathcal{U},u}$. To this end, let $\mathbf{V}\subseteq\mathbf{O}_{\mathcal{U},u}$ be open and nonempty (notice that $\mathcal{C}$ need not belong to $\mathbf{V}$). By Corollary 5.4, there are a brick partition and a walk $(\mathcal{V},v)\preceq(\mathcal{U},u)$ such that $\mathbf{O}_{\mathcal{V},v}\subseteq\mathbf{V}$. Up to a refinement, by (Upgrading Walks to Paths)., we can suppose that $v$ is a spaced path, and that $\mathcal{V}$ is fine enough that
@@
By picking $\mathcal{V}$ fine enough we can moreover assume there is $V_{0}\in\mathcal{V}(u(0))$ such that $x_{0}\in\operatorname{Cl}(V_{0})$ and such that $[V_{0}]_{1}\subseteq\mathcal{V}(u(0))$. Fix a path $P$ from $V_{0}$ to $v(0)$ in $\mathcal{V}(u(0))$. We will move $x_{0}$ along $P$, step by step. As a preliminary move, let $\mathcal{C}_{-1}\coloneqq\mathcal{C}$ and, using Proposition 7.5, find $g_{-1}\in G_{[V_{0}]_{1}}\subseteq G_{\mathcal{V},1}$ mapping $x_{0}$ to some $y_{0}\in V_{0}=P(0)$ and set $\mathcal{C}_{0}\coloneqq g_{-1}\cdot\mathcal{C}$. Note that $\mathcal{C}_{0}\in\mathbf{O}_{\mathcal{U},u}$ since $\mathcal{C}\in\mathbf{O}_{\mathcal{U},u}$ and $g_{-1}$ fixes every element of $u$ setwise. Suppose next we are given $\mathcal{C}_{i}\in\mathbf{O}_{\mathcal{U},u}$ with root $y_{i}\in P(i)$, for some $i<\operatorname{lh}(P)-1$. By Proposition 7.5, there is $g_{i}\in G_{P(i)\vee P(i+1)}$ which maps $y_{i}$ to some point $y_{i+1}\in P(i+1)$. Again, since $g_{i}$ fixes every element of $u$ setwise, $\mathcal{C}_{i+1}\coloneqq g_{i}\cdot\mathcal{C}_{i}$ belongs to $\mathbf{O}_{\mathcal{U},u}$.
@@
Suppose that $\mathcal{D}_{j}\in\mathbf{O}_{\mathcal{U},u}$ with root $z_{0}$ is given, and $\mathcal{D}_{j}\in\mathbf{O}_{\mathcal{V},\langle v(0),\dots,v(j)\rangle}$ for some $j<\operatorname{lh}(v)-1$. Let $D_{j}\in\mathcal{D}_{j}$ be such that $D_{j}\in O(v(0),v(1),\dots,v(j))$, and let $t_{j}\in v(j)\cap D_{j}$. In case $j=0$, make sure that $t_{0}\not=z_{0}$. By Proposition 7.5, there is $h_{j}\in G_{v(j)\vee v(j+1)}$ which maps $t_{j}$ to some point $t_{j+1}\in v(j+1)$. Again, if $j=0$, pick $h_{j}$ so that it also satisfies $h_{0}(z_{0})=z_{0}$. Call $\mathcal{D}_{j+1}\coloneqq h_{j}\cdot\mathcal{D}_{j}$. Notice that the root of $\mathcal{D}_{j+1}$ is $h_{j}(z_{0})=z_{0}$, since for $j>0$ the set $v(0)$ is pointwise fixed by $h_{j}$. Moreover, $h_{j}[D_{j}]\subseteq\bigvee_{k\leq j+1}v(k)$, as the set on the right is setwise fixed by $h_{j}$. Lemma 7.2 then implies
@@
It is left to prove that $\mathcal{D}_{j+1}\in\mathbf{O}_{\mathcal{U},u}$. If $\rho^{\mathcal{V}}_{\mathcal{U}}\cdot\langle v(0),\dots,v(j+1)\rangle\sqsubseteq u$, then $\mathbf{O}_{\mathcal{V},\langle v(0),\dots,v(j+1)\rangle}\subseteq\mathbf{O}_{\mathcal{U},u}$, in which case we conclude by (7.4). Otherwise, it holds that $v(j)\vee v(j+1)\subseteq\bigvee_{k<\operatorname{lh}(u)-1}u(k)$. By inductive assumption $\mathcal{D}_{j}\in\mathbf{O}_{\mathcal{U},u}$, so there is $A\in\mathcal{D}_{j}$ with $A\in O(u(0),\dots,u(-1))$. Therefore $A\cap u(-1)\neq\emptyset$, and since $h_{j}$ fixes $u(-1)$, we can infer $h_{j}[A]\cap u(-1)\neq\emptyset$. Moreover, $h_{j}[A]\subseteq\bigvee_{k<\operatorname{lh}(u)}u(k)$ since the set on the right is setwise fixed by $h_{j}$. We conclude by Lemma 7.2.
@@
Let $X$ be a strongly locally homogeneous Peano continuum which is not the circle, the sphere, or the real projective plane. Then $\operatorname{Homeo}(X)$ does not have the generic point property. In particular, its universal minimal flow is not metrizable and has no comeager orbit.